OpenStudy (anonymous):

solve using the quadratic formula y^2+9y+9=0

OpenStudy (mathmale):

this equation resembles ax^2 + bx + c = 0, the standard form of a quadratic equation. Compare this to your y^2 + 9y + 9 = 0. What are your coefficients a, b and c?

OpenStudy (anonymous):

do you do 9 divded by 2 and then square by 2

OpenStudy (anonymous):

but i dont get how to do it with fractions

OpenStudy (tkhunny):

If you were Completing the Square, that would be the right idea. If you're using the Quadratic Formula, according to your instructions, then no. Use the Quadratic Formula. Look at mathmale's response very carefully.

OpenStudy (mathmale):

Your y^2 + 9y + 9 = 0 could be re-written as 1y^2 + 9y + 9 = 0. Look at the three coefficients of the powers of y: They are 1, 9 and 9. Thus, your a =1 Your b = 9; and Your c = 9. Use the quadratic formula to solve for the two values of y that satisfy this equation.

OpenStudy (anonymous):

you know factorization method if you do apply that

OpenStudy (anonymous):

it will be easy and can`t take long time

OpenStudy (anonymous):

but i think his problem is telling him to use the quadratic formula @arush

OpenStudy (anonymous):

But its not mentioned here if not we can apply any method to solve quadratic equation

OpenStudy (mathmale):

I disagree. The problem statement clearly says "using the quadratic formula."

OpenStudy (anonymous):

ohh sorry then you are right

OpenStudy (mathmale):

Please review the above. a=1, b=9 and c=9. @kellyking: Please substitute these values of a, b and c into the quadratic formula, \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Note that you will end up with TWO distinct roots / solutions.

OpenStudy (anonymous):

yes value will be real or not depends on the discriminant

OpenStudy (mathmale):

Let's wait for a response from the questioner, @kellyking.

OpenStudy (anonymous):

ya